3. The magnetic field of an electromagnetic wave is given as B-3uT sin( 1x10 x -2xfi) where r is in meters. a) Calculate wavelength and frequency. b) Calculate the energy of one photon, in Joules and in eV c) Calculate the electric field amplitude.
Answer #1

As per the equation of the magnetic field of electromagnetic wave

B_z = 3\mu T Sin(1*10^7x-2\pi ft)

standard equation of the magnetic field of electromagnetic wave is

B_z = B_0 Sin[2\pi((\frac{x}{\lambda }) +( \frac{t}{T}))]

on comparing given equation with standard equation we have

B_0 = 3*10^{-6} T

2\pi(\frac{x}{\lambda }) = 1*10^7x

\frac{2\pi}{1*10^7 }= \lambda

\lambda= 6.283*10^{-7} m

frequency of the wave can be calculated by

v = \frac{c}{\lambda}= \frac{3*10^{8}}{6.283*10^{-7}} Hz

v = 4.775*10^{14}} Hz

------------------------- Solution of part b ----------------

Energy of photon is given by

E = hv = 6.626*10^{-34}*4.775*10^{14}}

E = 3.164*10^{-19}} J or

E = \frac{3.164*10^{-19}} {1.602*10^{-19}} eV

E = 1.975 eV

----------------- Solution of part C ---------------------------

To calculate the electric field amplitude we have

E_0=B_0*c = 3*10^{-6}*3.0*10^8 V/m

E_0=9.0*10^2 V/m